OpenStudy (anonymous):

How would I solve using LCD? (x-3)/(x-4) +4 = (3x)/(x)

OpenStudy (anonymous):

\[\frac{ x-3 }{ x-4 }+4 = \frac{ 3x }{ x }\]

OpenStudy (anonymous):